# MedHallBench: A New Benchmark for Assessing Hallucination in Medical Large Language Models

### v2

@@ Introduction
Large language models (LLMs) have demonstrated unparalleled capabilities in interpreting complex medical texts, patient records, and clinical notes, significantly advancing the fields of diagnosis, treatment planning, and patient care (Tian et al. 2023). However, despite these substantial achievements, medical LLMs are not without challenges. A critical issue is the tendency of these models to generate ”hallucinations”—medically unreliable or inaccurate information(Zhao et al. 2023). This phenomenon is not merely a minor flaw; it poses a serious threat to the integrity and reliability of AI applications in medicine(Singhal et al. 2023). The emergence of hallucinations within these models severely undermines their perceived competence and trustworthiness, raising urgent concerns about their credibility and safe application in clinical settings(Singhal et al. 2023). These hallucinations not only compromise accuracy but can also lead to grave consequences, such as misdiagnoses or inappropriate treatment plans, which could potentially harm patients(Wang et al. 2024). Furthermore, while comprehensive benchmarks exist for LLMs in other domains, such as pure language models and multimodal models, there is a lack of standardized evaluation frameworks for medical LLMs. The development of a benchmark for medical LLMs is crucial for guiding future advancements in this field.
@@
The identification of discrepancies between LLM-generated content and expert medical descriptions forms a crucial component of our evaluation framework. Through systematic comparison analysis, we employ a rigorous methodology to detect variations between model outputs and established medical knowledge. This process involves detailed examination of terminology usage, diagnostic reasoning, treatment recommendations, and clinical interpretations, enabling precise pinpointing of areas where the model deviates from expert consensus. The identification process is supported by a structured review protocol that ensures consistent and thorough examination of all content elements. Our methodology incorporates a comprehensive classification system for hallucination types, categorizing them based on their nature, severity, and relevance to medical facts. This classification framework distinguishes between various forms of hallucination, including factual inconsistencies, logical contradictions, temporal discrepancies, and clinical reasoning errors. Each category is further subdivided based on potential clinical impact and relationship to established medical knowledge, allowing for nuanced analysis of how different types of hallucinations might affect medical decision-making processes. This detailed categorization enables more targeted approaches to model improvement and refinement. Documentation of hallucination instances follows a systematic protocol designed to capture both quantitative and qualitative aspects of each occurrence. We maintain detailed records that include the context in which hallucinations occur, their specific manifestations, and their potential implications for medical outcomes. This documentation process encompasses comprehensive metadata collection, including the clinical context, the type of medical information involved, and the potential severity of impact on patient care. Such detailed documentation not only facilitates subsequent analysis but also provides valuable insights for developing mitigation strategies and improving model performance in critical medical applications. The specific flowchart of benchmark design is shown as Figure2
@@
In order to systematically study the performance of different models on hallucination detection tasks in MedHallBench, we used multiple evaluation metrics. In addition to the two ACHMI indicators mentioned above, we also employed various traditional indicators, including BertScore, METEOR, ROUGE-1/2/L, and BLEU.We also compare with other base models, including BLIP2 (Li et al. 2023c), InstructBLIP-7b/13b (Dai et al. 2023), LLaVA1.5-7b/13b(Liu et al. 2023a), mPLUGOwl2
(Ye et al. 2023), XrayGPT
(Thawkar et al. 2023), MiniGPT4
(Zhu et al. 2023),
RadFM
(Wu et al. 2023).
and LLaVA-Med(Li et al. 2023a) after fine-tuning on the Slake (SF) dataset.

### v3

@@ Introduction
Large language models (LLMs) have demonstrated unparalleled capabilities in interpreting complex medical texts, patient records, and clinical notes, significantly advancing the fields of diagnosis, treatment planning, and patient care (Tian et al. 2023; Tang et al. ; Zuo et al. 2024). However, despite these substantial achievements, medical LLMs are not without challenges. A critical issue is the tendency of these models to generate ”hallucinations”—medically unreliable or inaccurate information(Zhao et al. 2023). This phenomenon is not merely a minor flaw; it poses a serious threat to the integrity and reliability of AI applications in medicine(Singhal et al. 2023). The emergence of hallucinations within these models severely undermines their perceived competence and trustworthiness, raising urgent concerns about their credibility and safe application in clinical settings(Singhal et al. 2023). These hallucinations not only compromise accuracy but can also lead to grave consequences, such as misdiagnoses or inappropriate treatment plans, which could potentially harm patients(Wang et al. 2024). Furthermore, while comprehensive benchmarks exist for LLMs in other domains, such as pure language models and multimodal models, there is a lack of standardized evaluation frameworks for medical LLMs. The development of a benchmark for medical LLMs is crucial for guiding future advancements in this field.
@@
The identification of discrepancies between LLM-generated content and expert medical descriptions forms a crucial component of our evaluation framework. Through systematic comparison analysis, we employ a rigorous methodology to detect variations between model outputs and established medical knowledge. This process involves detailed examination of terminology usage, diagnostic reasoning, treatment recommendations, and clinical interpretations, enabling precise pinpointing of areas where the model deviates from expert consensus. The identification process is supported by a structured review protocol that ensures consistent and thorough examination of all content elements.
Our methodology incorporates a comprehensive classification system for hallucination types, categorizing them based on their nature, severity, and relevance to medical facts. This classification framework distinguishes between various forms of hallucination, including factual inconsistencies, logical contradictions, temporal discrepancies, and clinical reasoning errors. Each category is further subdivided based on potential clinical impact and relationship to established medical knowledge, allowing for nuanced analysis of how different types of hallucinations might affect medical decision-making processes. This detailed categorization enables more targeted approaches to model improvement and refinement.
Documentation of hallucination instances follows a systematic protocol designed to capture both quantitative and qualitative aspects of each occurrence. We maintain detailed records that include the context in which hallucinations occur, their specific manifestations, and their potential implications for medical outcomes. This documentation process encompasses comprehensive metadata collection, including the clinical context, the type of medical information involved, and the potential severity of impact on patient care. Such detailed documentation not only facilitates subsequent analysis but also provides valuable insights for developing mitigation strategies and improving model performance in critical medical applications.
@@
To systematically evaluate the performance of various models on hallucination detection tasks within MedHallBench, we employed multiple evaluation metrics. In addition to the two ACHMI indicators previously mentioned, we utilized traditional metrics such as BertScore, METEOR, ROUGE-1/2/L, and BLEU. We also compared the following base models: BLIP2 (Li et al. 2023c), InstructBLIP-7b/13b (Dai et al. 2023), LLaVA1.5-7b/13b (Liu et al. 2023a), mPLUGOwl2 (Ye et al. 2023), XrayGPT (Thawkar et al. 2023), MiniGPT4 (Zhu et al. 2023), RadFM (Wu et al. 2023), and LLaVA-Med (Li et al. 2023a). All the above models were fine-tuned on the Slake (SF) dataset.